# Hook-decomposable modules and their resolutions

### 24 Mar 2026 v1

@@ Abstract
We compare several classes of biparameter persistence modules: $\gamma$-products of monoparameter modules, hook-decomposable modules, modules admitting a Smith-type structure theorem, and modules of projective dimension 1. We determine all logical implications among these classes, providing explicit counterexamples showing that the converses fail when appropriate. In particular, $\gamma$-products (i.e., hook-decomposable modules) form a very small subclass of biparameter modules, precisely the ones for which a structure theorem still holds, thus making explicit the richer structural complexity of the biparameter setting compared to the monoparameter one.
@@ 1 Motivation and contribution
In previous work [10], we introduced one such construction, the $\gamma$-product, and showed that the resulting biparameter modules coincide precisely with hook-decomposable modules. Motivated by this, in the present work we analyze the logical relations between four classes of biparameter persistence modules naturally arising from, or closely related to, this construction: $(i)$ products of monoparameter modules (i.e., $\gamma$-products), $(ii)$ hook-decomposable modules, $(iii)$ modules for which a Smith-type structure theorem applies, $(iv)$ modules of projective dimension 1, $(v)$ modules of projective dimension 1 whose minimal free resolution splits. We show that $(i)$, $(ii)$, and $(iii)$ are equivalent, and that $(v)$ implies $(ii)$, which in turn implies $(iv)$; none of these last two implications has a converse. The resulting diagram of implications is given in Figure 1, while a schematic arrangement of these classes, as it will follow from the implications proved below, is shown in Figure 2.
@@ 2 Background notions
    Projective dimension. The module $M$ has projective dimension 1 if and only if $\beta_{2}=0$ in its minimal graded free resolution.
  6. 6.
    Split exact sequence. A short exact sequence $0\to A\to B\to C\to 0$ splits if $B\cong A\oplus C$, equivalently if the injection admits a retraction or the surjection admits a section.
@@ 3 Hook-decomposability: sufficient and necessary conditions
    If $M$ has projective dimension 1 and split resolution, then $M$ is hook-decomposable.
@@
    If $M$ is hook-decomposable, then $M$ has projective dimension 11.
@@
    From the hypothesis it follows that $M$ is composed by null or free summands, hence hook summands.
  2. 2.
    From Theorem 3.1, it follows that each hook summand is of the form $\mathbb{F}[x,y](-\mathbf{p})$ or $\big(\mathbb{F}[x,y]\big/(\mathbf{x}^{\,\mathbf{q}-\mathbf{p}})\big)(-\mathbf{p})$ and then their resolutions are $0\rightarrow\mathbb{F}[x,y](-\mathbf{p})\rightarrow\mathbb{F}[x,y](-\mathbf{p})\rightarrow 0$ and $0\rightarrow\mathbb{F}[x,y](-\mathbf{q})\xrightarrow{\;\mathbf{x}^{\,\mathbf{q}-\mathbf{p}}\;}\mathbb{F}[x,y](-\mathbf{p})\rightarrow\big(\mathbb{F}[x,y]\big/(\mathbf{x}^{\,\mathbf{q}-\mathbf{p}})\big)(-\mathbf{p})\rightarrow 0,$ respectively.
@@
    The module with one generator $g$ in bidegree $(0,0)$ and the relation xygxyg in bidegree $(1,1)$ is hook-decomposable and has projective dimension 1, yet its resolution does not split , since the injection admits no linear retraction.

### 15 Apr 2026 v2

@@ Abstract
We compare several classes of biparameter persistence modules: $\gamma$-products of monoparameter modules, hook-decomposable modules, modules admitting a Smith-type structure theorem, and modules of projective dimension at most 1. We determine all logical implications among these classes, providing explicit counterexamples showing that the converses fail when appropriate. In particular, $\gamma$-products (i.e., hook-decomposable modules) form a very small subclass of biparameter modules, precisely the ones for which a structure theorem still holds, thus making explicit the richer structural complexity of the biparameter setting compared to the monoparameter one.
@@ 1 Motivation and contribution
In previous work [10], we introduced one such construction, the $\gamma$-product, and showed that the resulting biparameter modules coincide precisely with hook-decomposable modules. Motivated by this, in the present work we analyze the logical relations between four classes of biparameter persistence modules naturally arising from, or closely related to, this construction: $(i)$ products of monoparameter modules (i.e., $\gamma$-products), $(ii)$ hook-decomposable modules, $(iii)$ modules for which a Smith-type structure theorem applies, $(iv)$ modules of projective dimension 1, $(v)$ free modules. We show that $(i)$, $(ii)$, and $(iii)$ are equivalent, and that $(v)$ implies $(ii)$, which in turn implies $(iv)$; none of these last two implications has a converse. The resulting diagram of implications is given in Figure 1, while a schematic arrangement of these classes, as it will follow from the implications proved below, is shown in Figure 2.
@@ 2 Background notions
    Projective dimension. The module $M$ has projective dimension $\leq 1$ if and only if $\beta_{2}=0$ in its minimal free resolution.
Remark $2.$
Please note that freeness (condition $(v)$) is equivalent to projective dimension $\leq 0$ and to the minimal free graded resolution being split short exact.
@@ 3 Hook-decomposability: sufficient and necessary conditions
    $(v)\Rightarrow(ii)$. If $M$ is free, then $M$ is hook-decomposable.
@@
    $(ii)\Rightarrow(iv)$. If $M$ is hook-decomposable, then $M$ has projective dimension $\leq 1$.
@@
  1. Each free summand of $M$ is either null or has a relation at infinity.
@@
    From Theorem 3.1, it follows that each hook summand is of the form $\mathbb{F}[x,y](-\mathbf{p})$ or
    $\big(\mathbb{F}[x,y]\big/(\mathbf{x}^{\,\mathbf{q}-\mathbf{p}})\big)(-\mathbf{p})$, so their resolutions are $0\rightarrow\mathbb{F}[x,y](-\mathbf{p})\rightarrow\mathbb{F}[x,y](-\mathbf{p})\rightarrow 0$ and $0\rightarrow\mathbb{F}[x,y](-\mathbf{q})\xrightarrow{\;\mathbf{x}^{\,\mathbf{q}-\mathbf{p}}\;}\mathbb{F}[x,y](-\mathbf{p})\rightarrow\big(\mathbb{F}[x,y]\big/(\mathbf{x}^{\,\mathbf{q}-\mathbf{p}})\big)(-\mathbf{p})\rightarrow 0,$ respectively.
@@
    The module with one generator $g$ in bidegree $(0,0)$ and the relation xygxyg in bidegree $(1,1)$ is hook-decomposable yet its not free. Notably, its resolution does not split, as the injection admits no graded retraction.